the length of a rectangle is decreased by 3 feet, and the width is increased by 1 foot, forming a square region havig an area of 25 square feet. what is the area of the orginal rectangle?
Answer by stanbon(75887)

the length of a rectangle is decreased by 3 feet, and the width is increased by 1 foot, forming a square region havig an area of 25 square feet. what is the area of the orginal rectangle?
Original Rectangle:
Length = x; Width=y
Area = xy
New SQUARE:
Length x-3; Width y+1
Area = 25 sq. ft.
Since it is a square and the area is 25 sq ft,
each side is 5 ft.
Therefore:
x-3=5 so x=8 ft
y+1=5 so y=4 ft
Therefore Area of the rectangle = xy=8*4= 32 sq ft.
Cheers,
Stan H.